OpenStudy (anonymous):

simplify. x^2-4x-21 over x-3 times x^2-5x+6 over x-2

OpenStudy (anonymous):

x^2 -4x - 21 = x^2 -7x + 3x -21 = x(x-7) +3(x-7) = (x+3)(x-7) x^2 -5x +6 = x^2 -3x -2x +6 = x(x-3) -2(x-3) = (x-2)(x-3) ((x+3)(x-7)/(x-3)) * ((x-2)(x-3)/(x-2)) ((x+3)(x-7)(x-2)(x-3))/((x-3)(x-2)) = (x+3)(x-7)

OpenStudy (anonymous):

thankyou!